# Haitao Xu Corresponding author    Jingru Zhang

### v1

@@ Abstract
In this paper, we study the $k$-center problem of uncertain points on a graph. Given are an undirected graph $G=(V,E)$ and a set $\mathcal{P}$ of $n$ uncertain points where each uncertain point with a non-negative weight has $m$ possible locations on $G$ each associated with a probability. The problem aims to find $k$ centers (points) on $G$ so as to minimize the maximum weighted expected distance of uncertain points to their expected closest centers. No previous work exist for the $k$-center problem of uncertain points on undirected graphs. We propose exact algorithms that solve respectively the case of $k=2$ in $O(|E|^{2}m^{2}n\log|E|mn\log mn)$ time and the problem with $k\geq 3$ in $O(\min\{|E|^{k}m^{k}n^{k+1}k\log|E|mn\log m,|E|^{k}n^{\frac{k}{2}}m^{\frac{k^{2}}{2}}\log|E|mn\})$ time, provided with the distance matrix of $G$. In addition, an $O(|E|mn\log mn)$-time algorithmic approach is given for the one-center case.
Keywords:
[…]
@@ 1 Introduction
Facility location problems on uncertain demands have drawn growing attentions [10, 15, 23, 21, 1, 2] in academia due to the inherent uncertainty of collected demand data. In this paper, we study the network $k$-center problem, a classical problem in facility locations, for uncertain points (demands) where each uncertain point models a customer with multiple possible locations on the network.
Let $G=(V,E)$ be the given undirected graph. For any two vertices $u,v$ of $G$, denote by $e(u,v)$ the edge of length $\overline{e(u,v)}>0$ connecting $u$ and $v$. Any point $p$ on $e(u,v)$ is represented by a tuple $(u,v,t(p))$ where $t(p)$ is the distance (the path length) of $p$ to $u$ along $e(u,v)$. For any two points $p$ and $q$ of $G$, the distance $d(p,q)$ between them is the length of their shortest path $\pi(p,q)$.
Let $\mathcal{P}$ be a set of $n$ uncertain points $P_{1},\cdots,P_{n}$. Each $P_{i}$ is associated with $O(m)$ points (locations) $p_{i1},\cdots,p_{im}$ on $G$ where each $p_{ij}$ is attached with a value $f_{ij}\geq 0$ so that $P_{i}$ appears at $p_{ij}$ with the probability $f_{ij}$. Additionally, each $P_{i}\in\mathcal{P}$ is of weight $w_{i}\geq 0$. For any (certain) point $p$ of $G$, the distance of $P_{i}$ to $p$ is the expected version and defined as $\sum_{j=1}^{m}f_{ij}\cdot d(p_{ij},p)$, denoted by $\mathsf{E}\mathrm{d}(P_{i},p)$.
The $k$-center problem is to find a set $Q$ of $k$ points (centers) in $G$ so that the maximum weighted expected distance of uncertain points in $\mathcal{P}$ to their expected closest points in $Q$ is minimized. Namely, it aims to find a set $Q$ of $k$ points in $Q$ so as to minimize the objective value $\phi(\mathcal{P},Q)=\max_{P_{i}\in\mathcal{P}}\min_{q\in Q}\{w_{i}\mathsf{E}\mathrm{d}(P_{i},q)\}$.
When $G$ is a path, the problem can be addressed in $O(mn\log mn+n\log k\log n)$ time [20]. For $G$ being a tree graph, this problem can be solved in $O(n^{2}\log n\log mn+mn\log^{2}mn\log n)$ time [21]. On a general graph, the problem is NP-hard due to the NP-hardness of the case $m=1$. To the best of our knowledge, no known algorithms exist for solving it. In this paper, we give the first exact algorithms for solving the problem respectively under $k=2$ and $k\geq 3$, and develop a faster algorithm for the case of $k=1$.
@@
For certain point, i.e., $m=1$, the network $k$-center problem has been explored a lot in the literature. Kariv and Hakimi [16] proposed the first exact algorithm that solves the problem on a vertex-weighted graph in $O((|E|^{k}\cdot n^{2k-1}/(k-1)!)\cdot\lg n)$. Later, Tamir [19] gave a faster exact algorithm that solves the problem in $O(m^{k}n^{k}\log^{2}n)$ time. The state-of-art result was given by Bhattacharya and Shi [7] and their algorithm computes the exact solution in $O(m^{k}n^{k/2}2^{\log^{*}n}\log n)$ time by remodeling the decision problem to a Klee’s measure problem which asks for the measure of the union of axis-parallel boxes. Moreover, faster exact algorithms [18, 17, 4, 3, 6, 12, 22] were proposed for the problem on special graphs including tree and cactus graphs or the problem with a constant $k$.
When it comes to uncertain points each with multiple possible locations, as the above mentioned, the $k$-center problem was studied first on the path network by Wang and Zhang [20]. Later, they studied the problem on tree networks and gave an $O(n^{2}\log n\log mn+mn\log^{2}mn\log n)$-time algorithm [21]. In addition, the problem was considered for a constant $k$ respectively on a path, a tree, and a cactus graph, and faster algorithms were developed [14, 24, 25, 26]. On general graphs, as far as we are aware, no previous work exist for the problem even with $k=1$. Also, see the algorithm works [20, 23, 14] for the path $k$-center problem of uncertain points with appearance of line segments.
@@ 2 Preliminaries
Let $p$ be any point of $G$. Provide with the distance matrix, for any $P_{i}\in\mathcal{P}$, value $\mathsf{E}\mathrm{d}(P_{i},p)$ can be computed in $O(m)$ time. Because the shortest path between any two points can be obtained in constant time.
Let $e$ be an edge of $G$ that is incident to vertices $u$ and $v$ of $G$. Consider function $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e$. Recall that $x=(u,v,t(x))$ is a point of $e$ at distance $t(x)$ to $u$ along $e$. We have the following lemma.
@@
$\mathsf{E}\mathrm{d}(P_{i},x)$ is a piecewise linear function of complexity $O(m)$ and it can be determined in $O(m\log m)$ time.
@@
Consider the distance $d(p,x)$ between point $p$ and $x$. On the one hand, $p$ is not in $e/\{u,v\}$, i.e., $p^{\prime}$ is in the interior of $e$. As illustrated in Fig.2 (a), as $x$ moves from $u$ to $v$ along $e$, $d(p,x)$ increases, decreases, or first increases until $x$ reaches the semicircular point $p^{\prime}$ of $p$ on $e$, which is defined later, and then starts to decrease [16]. If $d(p,u)+\overline{e}>d(p,v)$ and $d(p,v)+\overline{e}>d(p,u)$, then there exists a (unique) point $p^{\prime}$ in the interior of $e$ so that $d(p,u)+t(p^{\prime})=d(p,v)+\overline{e}-t(p^{\prime})$ at $x=p^{\prime}$, that is, $d(p,x)$ falls into the third case; $p^{\prime}$ is referred to as $p$’s semicircular point on $e$. For the first two cases, no semicircular points exist for $p$ on $e$.
In details, if $d(p,u)+\overline{(}e)=d(p,v)$ (i.e., the first case of $d(p,x)$), then $d(p,x)=d(p,u)+t(x)$ at any $x\in e$. If $d(p,u)-\overline{(}e)=d(p,v)$ (i.e., the second case of $d(p,x)$), $d(p,x)=d(p,v)+\overline{e}-t(x)$ at any $x\in e$. Otherwise, $d(p,x)=d(p,u)+t(x)$ for any $x$ in the segment $[u,p^{\prime}]$ on $e$, and $d(p,x)=d(p,v)+\overline{e}-t(x)$ for $x\in[p^{\prime},v]$ where $p^{\prime}$ can be found in $O(1)$ time.
Figure 1: Illustrating function $d(p,x)$ of a point $p$ for $x$ in edge $e(u,v)$. (a) is for the case where $p$ is not in $e/\{u,v\}$ while (b) is for the situation where $p$ is in the interior of $e$.
@@
On the other hand, $p$ is in the interior of $e$, i.e., $p\in e/\{u,v\}$. As illustrated in Fig.2 (b), if $d(u,v)<\overline{e}$, then $p$ must have an unique point $p^{\prime}$ in the interior of $e$ so that $d(p,u)+t(p^{\prime})=t(p)-t(p^{\prime})=d(p,p^{\prime})$ where $d(p,u)=d(p,v)+d(u,v)$; we also say $p^{\prime}$ is $p$’s semicircular point on $e$, and $p^{\prime}$ can be known in $O(1)$ time. If so, $d(p,x)=t(x)+d(p,u)$ for any $x$ in segment $[u,p^{\prime}]$ on $e$, and $d(p,x)=|t(p)-t(x)|$ for any $x\in[p^{\prime},v]$ on $e$. Otherwise, $d(u,v)>\overline{e}$ so $d(p,x)=|t(p)-t(x)|$ for any $x\in e$.
@@
Thus, $d(p,x)$ for $t(x)\in e$ is a piecewise linear function of $O(1)$ complexity; it turns at $p$’s semicircular points $p^{\prime}$ and $p$ if $p^{\prime}$ exists and $p$ is in the interior of $e$. Refer to any point on $e$ where $d(p,x)$ turns as its turning point. It follows that $\mathsf{E}\mathrm{d}(P_{i},x)$ is a piecewise linear function which turns at each turning point of functions $d(p_{ij},x)$ for all $1\leq j\leq m$, and its complexity is $O(m)$.
@@
Now we discuss how to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e$. First, we determine $d(p_{ij},x)$ in $x\in e$ for all $1\leq j\leq m$, and then sort all their turning points on $e$ by their distances to $u$ along $e$, which can be carried out totally in $O(m\log m)$ time. Join $u$ and $v$ to this ordered set. We continue to scan the set to remove duplicates while keep for each distinct turning point a list that contains all locations of $P_{i}$ whose distance functions $d(p_{ij},x)$ turn at this turning point. Let $X_{i}=\{x_{i1},\cdots,x_{i,m_{i}}\}$ be the obtained (ordered) set of $\mathsf{E}\mathrm{d}(P_{i},x)$’s turning points on $e$, where $m_{i}=O(m)$, $x_{i1}=u$, and $x_{i,m_{i}}=v$; denote by $L_{i,s}$ the computed for $x_{i,s}$ that contains all locations whose functions $d(p_{ij},x)$ turn at $x_{i,s}$. Set $X_{i}$ can be obtained totally in $O(m\log m)$ time.
@@
Proceed to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ over each segment $[x_{i,s},x_{i,s+1}]$ for all $11\leq s\leq m_{i}-$ 1 in order. Loop $X_{i}$ from the beginning. For $x_{i1}$, we sum $f_{ij}d(p_{ij},x)$ at $x=u$ for all $1\leq j\leq m$ to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ over $[x_{i1},x_{i2}]$, and attach the obtained function with $x_{i1}$ in $X_{i}$. For each $x_{i,s}$ with $2\leq s<m_{i}$, letting $g(x)$ represent $\mathsf{E}\mathrm{d}(P_{i},x)$ over $[x_{i,s-1},x_{i,s}]$ (which is a linear function), we loop through the list $L_{i,s}$ of $x_{i,s}$ to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ over $[x_{i,s},x_{i,s+1}]$ as follows. For each location $p_{ij}$ in $L_{i,s}$, subtract $f_{ij}d(p_{ij},x)$ for $x\in[x_{i,s-1},x_{i,s}]$ from $g(x)$ and then add $f_{ij}d(p_{ij},x)$ for $x\in[x_{i,s},x_{i,s+1}]$ to $g(x)$. Because given any point $p$ of $e$, $d(p_{ij},x)$ at $x=p$ can be determined in $O(1)$ time. It takes $O(|L_{i,s}|)$ time to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in[x_{i,s},x_{i,s+1}]$. Due to $\sum_{s=1}^{s=m_{i}}L_{i,s}=O(m)$, it takes $O(m\log m)$ to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e$, provided with the distance matrix.
@@
Suppose that $\lambda^{*}$ falls into neither of the two cases. Let $x^{\prime}$ be a center that determines $\lambda^{*}$, and let $\mathcal{P}^{\prime}$ be the subset of uncertain points assigned to $x^{\prime}$. If $x^{\prime}$ is in the interior of an edge, e.g., $e(u,v)$, then one can move $x^{\prime}$ along $e$ toward one vertex to reduce the maximum weighted expected distance $\phi(\mathcal{P}^{\prime},x^{\prime})$ of $x^{\prime}$ to $\mathcal{P}^{\prime}$ until $x^{\prime}$ reaches a point $x^{\prime\prime}$ so that $(x^{\prime\prime},\phi(\mathcal{P}^{\prime},x^{\prime\prime}))$ is either an intersection of the weighted expected distance functions of two uncertain points in $\mathcal{P}^{\prime}$ where their slopes are of opposite signs, or a breakpoint of the weighted expected distance function of an uncertain point in $\mathcal{P}^{\prime}$, which leads a contradiction with the assumption. Thus, the lemma holds. ∎

### v2

@@ Abstract
In this paper, we study the $k$-center problem of uncertain points on a graph. Given are an undirected simple graph $G=(V,E)$ and a set $\mathcal{P}$ of $n$ uncertain points where each uncertain point with a non-negative weight has $m$ possible locations on $G$ each associated with a probability. The problem aims to find $k$ centers (points) on $G$ so as to minimize the maximum weighted expected distance of uncertain points to their own expected closest centers. No previous work exist for the $k$-center problem of uncertain points on undirected graphs. We propose exact algorithms that solve respectively the case of $k=2$ in $O(|E|^{2}m^{2}n\log|E|mn\log mn)$ time and the problem with $k\geq 3$ in $O(\min\{|E|^{k}m^{k}n^{k+1}k\log|E|mn\log m,|E|^{k}n^{\frac{k}{2}}m^{\frac{k^{2}}{2}}\log|E|mn\})$ time, provided with the distance matrix of $G$. In addition, an $O(|E|mn\log mn)$-time algorithmic approach is given for the one-center case.
@@ 1 Introduction
Facility location problems on uncertain demands have drawn growing attentions [10, 15, 23, 21, 1, 2] in academia due to the inherent uncertainty of collected demand data. In this paper, we study the network $k$-center problem, a classical problem in facility locations, for uncertain points where each uncertain point models a demand customer with multiple possible locations on the network.
Let $G=(V,E)$ be an undirected graph which contains no loops and no multiple edges, and where each edge is of a positive length. For any two vertices $u,v$ of $G$, denote their edge by $e(u,v)$ and the length of $e(u,v)$ by $\overline{e(u,v)}$. Assume $G$ is embedded in the Euclidean plane, so we can talk about “points” on an edge: A (deterministic) point $p$ on $e(u,v)$ is represented by a tuple $(u,v,t(p))$, where $t(p)$ is the distance of $p$ to $u$ along $e(u,v)$. For any two points $p$ and $q$ of $G$, the distance $d(p,q)$ between them is the sum of the lengths of all edges on their shortest path $\pi(p,q)$ along $G$.
An uncertain point of $G$ is a point whose location in $G$ is not certain and follows a discrete probability density function. Let $\mathcal{P}$ be a set of $n$ uncertain points $P_{1},\cdots,P_{n}$ in $G$. Each $P_{i}$ is associated with $m$ points $p_{i1},\cdots,p_{im}$ on $G$, where each $p_{ij}$ is attached with a value $f_{ij}\geq 0$ so that $P_{i}$ appears at point $p_{ij}$ with the probability $f_{ij}$. Additionally, each $P_{i}\in\mathcal{P}$ is of weight $w_{i}\geq 0$. For any point $p$ of $G$, the distance of $P_{i}$ to $p$ is defined as $\sum_{j=1}^{m}f_{ij}\cdot d(p_{ij},p)$, denoted by $\mathsf{E}\mathrm{d}(P_{i},p)$, which is the expected version of the distance between $P_{i}$ and $p$.
The $k$-center problem is to find a set $Q$ of $k$ points (centers) in $G$ so that the maximum weighted expected distance of uncertain points in $\mathcal{P}$ to their own expected closest points in $Q$ is minimized. Namely, it aims to find a set $Q$ of $k$ points in $Q$ so as to minimize the objective value $\phi(\mathcal{P},Q)=\max_{P_{i}\in\mathcal{P}}\min_{q\in Q}\{w_{i}\mathsf{E}\mathrm{d}(P_{i},q)\}$.
When $G$ is a path, the problem can be addressed in $O(mn\log mn+n\log k\log n)$ time [20]. For $G$ being a tree graph, this problem can be solved in $O(n^{2}\log n\log mn+mn\log^{2}mn\log n)$ time [21]. On a general graph, the problem is NP-hard due to the NP-hardness of the case $m=1$ [16, 18]. To the best of our knowledge, no known algorithms exist for solving it. In this paper, we give the first exact algorithms for solving the problem respectively under $k=2$ and $k\geq 3$, and develop a faster algorithm for the case of $k=1$.
@@
For deterministic points, i.e., $m=1$, the network $k$-center problem has been explored a lot in the literature. Kariv and Hakimi [16] proposed the first exact algorithm that solves the problem on a vertex-weighted graph in $O((|E|^{k}\cdot n^{2k-1}/(k-1)!)\cdot\lg n)$. Later, Tamir [19] gave a faster exact algorithm that solves the problem in $O(m^{k}n^{k}\alpha(n)\log^{2}n)$ time where $\alpha(n)$ is the inverse Ackermann function. The state-of-art result was given by Bhattacharya and Shi [7], and their algorithm computes the exact solution in $O(m^{k}n^{k/2}2^{\log^{*}n}\log n)$ time by remodeling the decision problem to a Klee’s measure problem, which asks for the measure of the union of axis-parallel boxes. Moreover, faster exact algorithms [18, 17, 4, 3, 6, 12, 22] were proposed for the problem on special graphs including tree and cactus graphs or the problem with a constant $k$.
When it comes to uncertain points each with multiple possible locations, as the above mentioned, the $k$-center problem was studied first on the path network by Wang and Zhang [20]. Later, they studied the problem on tree networks and gave an $O(n^{2}\log n\log mn+mn\log^{2}mn\log n)$-time algorithm [21]. In addition, the problem was considered for a constant $k$ respectively on a path, a tree, and a cactus graph, and faster algorithms were developed [14, 24, 25, 26]. On general graphs, as far as we are aware, no known work exists for the problem even with $k=1$. Also, see the algorithm works [20, 23, 14] for the path $k$-center problem of uncertain points with appearance of line segments.
@@ 2 Preliminaries
Since the problem on a tree network has been addressed in previous work, we consider only the general situation where $G$ is an undirected simple graph but not a tree. Hence, $|V|=O(|E|)$.
Provided with the distance matrix, for any two point $p,q$ of $G$, their distance $d(p,q)$ can be obtained in $O(1)$ time since the shortest path length between any two vertices of the edges containing $p$ and $q$ can be known in $O(1)$ time by the distance matrix. As a result, for any $P_{i}\in\mathcal{P}$, the expected distance $\mathsf{E}\mathrm{d}(P_{i},p)$ of $P_{i}$ to point $p$ can be computed in $O(m)$ time.
Recall that $e(u,v)$ is an arbitrary edge of $G$ which is incident to vertex $u$ and $v$. We abuse $e(u,v)$ to the set of all points on it. We say that a point is in the interior of edge $e(u,v)$ if it is in set $e(u,v)\setminus\{u,v\}$. For point $p$ and edge $e(u,v)$, we say that a point of $e(u,v)$ is a semicircular point of $p$ if this point has two shortest paths to $p$ such that the two paths or their subpaths form a cycle of $G$ including $e(u,v)$.
Recall that $x=(u,v,t(x))$ is a point of $e(u,v)$ at distance $t(x)$ to $u$ along $e(u,v)$. For any $P_{i}\in\mathcal{P}$, $\mathsf{E}\mathrm{d}(P_{i},x)$ is a function in $t(x)$ as $x$ moves along $e(u,v)$ from $u$ to $v$. We have the following lemma for $\mathsf{E}\mathrm{d}(P_{i},x)$.
@@
For any $P_{i}\in\mathcal{P}$, function $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e(u,v)$ is a piecewise linear function consisting of $O(m)$ pieces, and it can be explicitly computed in $O(m\log m)$ time.
@@
Let $p$ be any point of $G$. If $p\in e(u,v)$, suppose $p$ is at distance $t(p)$ to $u$ along $e(u,v)$. We first consider the distance $d(p,x)$ from $p$ to $x$.
On the one hand, $p$ is not in set $e(u,v)\setminus\{u,v\}$. As illustrated in Fig.2 (a), as $x$ moves from $u$ to $v$ along $e(u,v)$, $d(p,x)$ increases, decreases, or first increases until $x$ reaches a point of $e(u,v)$ and then starts to decrease [16]. In details, if $d(p,u)+\overline{e(u,v)}=d(p,v)$ (i.e., the first case of $d(p,x)$), then $d(p,x)=d(p,u)+t(x)$ at any $x\in e(u,v)$. If $d(p,v)+\overline{e(u,v)}=d(p,u)$ (i.e., the second case of $d(p,x)$), $d(p,x)=d(p,v)+\overline{e(u,v)}-t(x)$ at any $x\in e(u,v)$. Otherwise, $d(p,u)+\overline{e(u,v)}>d(p,v)$ and $d(p,v)+\overline{e(u,v)}>d(p,u)$; so $p$ has an unique point $p^{\prime}=(u,v,t(p^{\prime}))$ in the interior of $e(u,v)$ such that $d(p,u)+t(p^{\prime})=d(p,v)+\overline{e(u,v)}-t(p^{\prime})$, but $d(p,u)+t(x)<d(p,v)+\overline{e(u,v)}-t(x)$ for any $x$ in the segment $u,p^{\prime})$ of $e(u,v)$ and $d(p,u)+t(x)>d(p,v)+\overline{e(u,v)}-t(x)$ for $x\in(p^{\prime},v]$. By the definition, $p^{\prime}$ is the semicircular point of $p$ on $e(u,v)$. In this case, $d(p,x)=d(p,u)+t(x)$ for $x\in[u,p^{\prime}]$, and $d(p,x)=d(p,v)+\overline{e(u,v)}-t(x)$ for $x\in[p^{\prime},v]$.
Figure 1: Illustrating function $d(p,x)$ of a point $p$ for $x$ in edge $e(u,v)$. (a) is for the case where $p$ is not in $e(u,v)\setminus\{u,v\}$ while (b) is for the situation where $p$ is in the interior of $e(u,v)$.
@@
On the other hand, $p$ is in the interior of $e(u,v)$. As illustrated in Fig.2 (b), if $d(p,u)<t(p)$, i.e, $\pi(p,u)\notin e(u,v)$, then there is an unique point $p^{\prime}$ in the interior of $e(u,v)$ so that $d(p,u)+t(p^{\prime})=t(p)-t(p^{\prime})$, but d(p,u)+t(x)<t(p)t(x)d(p,u)+t(x)<t(p)-t(x) for any $x\in[u,p^{\prime})$ and d(p,u)+t(x)>t(p)t(x)d(p,u)+t(x)>t(p)-t(x) for $x\in(p^{\prime},v]$. Clearly, $p^{\prime}$ is the semicircular point of $p$ on $e(u,v)$. Hence, if $d(p,u)<t(p)$, $d(p,x)=t(x)+d(p,u)$ for $x\in[u,p^{\prime}]$ and d(p,x)=|t(p)t(x)|d(p,x)=|t(p)-t(x)| for any $x\in[p^{\prime},v]$. Likewise, if $d(p,v)<\overline{e(u,v)}-t(p)$, then there is an unique point $p^{\prime\prime}=(u,v,t(p^{\prime\prime}))$ in $e(u,v)$, which is between $p$ and $v$, so that $t(p^{\prime\prime})-t(p)=d(p,v)+\overline{e(u,v)}-t(p^{\prime\prime})$, but $|t(x)-t(p)|<d(p,v)+\overline{e(u,v)}-t(x)$ for any $x\in[u,p^{\prime\prime})$ and $t(x)-t(p)>d(p,v)+\overline{e(u,v)}-t(x)$ for any $x\in(p^{\prime\prime},v]$. Point $p^{\prime\prime}$ is $p$’s semicircular point on $e(u,v)$. Otherwise, $d(p,u)=t(p)$ and $d(p,v)=\overline{e(u,v)}-t(p)$, so d(p,x)=|t(p)t(x)|d(p,x)=|t(p)-t(x)| for any $x\in e(u,v)$.
@@
Thus, $d(p,x)$ for $x\in e(u,v)$ is a piecewise linear function consisting of $O(1)$ pieces; it turns at $p$’s semicircular point and $p$ if $p$ has a semicircular point on $e(u,v)$ and $p\in e(u,v)$. Refer to any point on $e(u,v)$ where $d(p,x)$ turns as its turning point. It follows that $\mathsf{E}\mathrm{d}(P_{i},x)$ is a piecewise linear function, which turns at each turning point of functions $d(p_{ij},x)$ for all $1\leq j\leq m$, and its graph consists of $O(m)$ pieces.
@@
Now we discuss how to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e(u,v)$. First, we determine $d(p_{ij},x)$ in $x\in e(u,v)$ for all $1\leq j\leq m$ and all their turning points on $e(u,v)$ in $O(m)$ time. Then, sort all turning points on $e(u,v)$ by their distances to $u$ along $e(u,v)$ in $O(m\log m)$ time. Join $u$ and $v$ to the ordered set of turning points. We continue to scan the set to remove duplicates while compute for each distinct turning point, a list that contains all locations of $P_{i}$ whose distance functions $d(p_{ij},x)$ turn at this turning point. Let $X_{i}=\{x_{i,1},\cdots,x_{i,m_{i}}\}$ be the obtained (ordered) set of $\mathsf{E}\mathrm{d}(P_{i},x)$’s turning points on $e(u,v)$, where $m_{i}=O(m)$, $x_{i,1}=u$, and $x_{i,m_{i}}=v$. Denote by $L_{i,s}$ the obtained list for $x_{i,s}$ which contains all $P_{i}$’s locations whose functions $d(p_{ij},x)$ turn at $x_{i,s}$. Set $X_{i}$ can be obtained totally in $O(m\log m)$ time.
@@
Proceed to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ over each segment $x_{i,s},x_{i,s+1}]$ for all $1\leq s\leq m_{i}-1$ in order. Loop $X_{i}$ from the beginning. For $x_{i1}$, we sum $f_{ij}d(p_{ij},x)$ at $x=u$ for all $1\leq j\leq m$ to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ over $x_{i,1},x_{i,2}]$, and attach the obtained function with $x_{i,1}$ in $X_{i}$. For each $x_{i,s}$ with $2\leq s<m_{i}$, letting $g(x)$ represent $\mathsf{E}\mathrm{d}(P_{i},x)$ over $x_{i,s-1},x_{i,s}]$ (which is a linear function), we loop through the list $L_{i,s}$ of $x_{i,s}$ to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ over $x_{i,s},x_{i,s+1}]$ as follows. For each location $p_{ij}$ in $L_{i,s}$, subtract $f_{ij}d(p_{ij},x)$ for $x\in[x_{i,s-1},x_{i,s}]$ from $g(x)$ and then add $f_{ij}d(p_{ij},x)$ for $x\in[x_{i,s},x_{i,s+1}]$ to $g(x)$. Recall that $d(p_{ij},x)$ at any point can be determined in $O(1)$ time. It thus takes $O(|L_{i,s}|)$ time to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in[x_{i,s},x_{i,s+1}]$. Due to $\sum_{s=1}^{s=m_{i}}L_{i,s}=O(m)$, it takes $O(m\log m)$ to determine $\mathsf{E}\mathrm{d}(P_{i},x)$ for $x\in e(u,v)$, provided with the distance matrix.
@@
Suppose that $\lambda^{*}$ falls into neither of the two cases. Let $x^{\prime}$ be a center that determines $\lambda^{*}$, and let $\mathcal{P}^{\prime}$ be the subset of uncertain points assigned to $x^{\prime}$. If $x^{\prime}$ is in the interior of an edge, e.g., $e(u,v)$, then one can move $x^{\prime}$ along $e(u,v)$ toward one vertex to reduce the maximum weighted expected distance $\phi(\mathcal{P}^{\prime},x^{\prime})$ of $x^{\prime}$ to $\mathcal{P}^{\prime}$ until $x^{\prime}$ reaches a point $x^{\prime\prime}$ so that $(x^{\prime\prime},\phi(\mathcal{P}^{\prime},x^{\prime\prime}))$ is either an intersection of the weighted expected distance functions of two uncertain points in $\mathcal{P}^{\prime}$ where their slopes are of opposite signs, or a breakpoint of the weighted expected distance function of an uncertain point in $\mathcal{P}^{\prime}$, which leads a contradiction with the assumption. Thus, the lemma holds. ∎